OpenStudy (anonymous):

find two consecutive whole numbers that square root of 86 lies between

OpenStudy (solomonzelman):

Some prefect squares are: \(\large\color{black}{ \displaystyle 49,~~~64,~~~81,~~100,~~~121}\)

OpenStudy (solomonzelman):

some little hint. does it help ?

OpenStudy (anonymous):

I still don't understand what Im suppose to do?

OpenStudy (solomonzelman):

For example, `find two consecutive whole numbers that square root of 21 lies between` you know that \(\large\color{black}{ \displaystyle 4^2=16 }\) \(\large\color{black}{ \displaystyle 5^2=25 }\) that means that \(\large\color{black}{ \displaystyle \sqrt{21} }\) is somewhere between \(\large\color{black}{ \displaystyle \sqrt{16} }\) and \(\large\color{black}{ \displaystyle \sqrt{25} }\), and therefore your answer is that \(\large\color{black}{ \displaystyle \sqrt{21} }\) is between 3 and 4.